OpenStudy (fo14):

1. |x|+5=18 (1 point) A: 5 or -5 B: 13 or -13 (I think it's this one) C: 18 or -18 D: 23 or -23 2. |y+4|< 1 (1 point ) A: -5

OpenStudy (lathrom1234):

a is right

OpenStudy (elise_a18):

^

OpenStudy (fo14):

is it all correct

OpenStudy (elise_a18):

3 is correct

OpenStudy (fo14):

ok

OpenStudy (fo14):

wait so is 3 wrong or right

OpenStudy (jango_in_dtown):

2' 2. |y+4|< 1 implies -1<y+4<1 implies -5<y<-3

OpenStudy (jango_in_dtown):

4 will be C rest all are correct

OpenStudy (fo14):

ok

OpenStudy (fo14):

i just tried it and i was right about 4 ur answer came out wrong

OpenStudy (fo14):

4 was a

OpenStudy (jango_in_dtown):

|a|-(3/4)=(5/8) or, |a|=(3/4)+(5/8)=11/8 or, a=+11/8. -11/8 11/8= 1 3/8

OpenStudy (jango_in_dtown):

It cant be wrong